# A New Approach for Fourier Extension Based on Weighted Generalized Inversethanks: The research is partially supported by National Natural Science Foundation of China (Nos. 12261131494, RSF-NSFC 23-41-00002)

### v2

@@ 3.3 The influence of the parameter $T$ on calculation accuracy
In actual calculations, (27) tells us that when approximating functions of the same frequency, an increase in $T$ means that a larger $N$ is required, and a larger $\gamma$ corresponds to a larger $M$, which will affect the final calculation efficiency. This is not the focus of this paper. For a fast algorithm based on boundary intervals data, please refer to [21].
@@ 4 Numerical tests
First, we show the approximation results of the following three test functions:
$f_{1}(t)=t^{2},\quad f_{2}(t)=\exp(t),\quad f_{3}(t)=x\exp(\sin(5.4\pi t)+\cos(\sqrt{2}\pi t)).$
The frequencies of these three test functions are not high. Figure 1 shows the weighted projection ${{\bf Q}}_{\breve{\bf L},N}^{\epsilon}{\bf f}$ and the corresponding approximation error obtained when $p$ takes different values. It can be seen that the approximation errors of different $p$ values are close. From remark 2.1, we know that $p=0$ corresponds to the original method, so the calculation result is oscillating in the extended region. As the $p$ value increases, the solution becomes smoother in the extended region, effectively overcoming the oscillation.
Our next collection of test functions contain high frequency:
$f_{4}(t)=\text{erf}(100t),\quad f_{5}(t)=\cos\left(\frac{100}{1+25t^{2}}\right),\quad f_{6}(t)=\frac{1}{1.1-t^{2}}.$
Figure 6 shows the corresponding calculation results. It can be seen that the high-frequency information at the boundary has a slightly greater impact on the calculation, but the method is still effective. The weighted operator changes the oscillation characteristics of the calculation results in the extended region.
(a) ${{\bf Q}}_{\breve{\bf L},N}^{\epsilon}{\bf f}$, $f_{1}(t)$, $N=200$.
(b) approximation error
(c) ${{\bf Q}}_{\breve{\bf L},N}^{\epsilon}{\bf f}$, $f_{2}(t)$, $N=200$.
(d) approximation error
(e) ${{\bf Q}}_{\breve{\bf L},N}^{\epsilon}{\bf f}$, $f_{3}(t)$, $N=200$.
(f) approximation error
Figure 5: ${{\bf Q}}_{\breve{\bf L},N}^{\epsilon}{\bf f}$ and their approximation errors.
(a) ${{\bf Q}}_{\breve{\bf L},N}^{\epsilon}{\bf f}$, $f_{4}(t)$, $N=800$.
(b) approximation error
(c) ${{\bf Q}}_{\breve{\bf L},N}^{\epsilon}{\bf f}$, $f_{5}(t)$, $N=800$.
(d) approximation error
(e) ${{\bf Q}}_{\breve{\bf L},N}^{\epsilon}{\bf f}$, $f_{6}(t)$, $N=800$.
(f) approximation error
Figure 6: ${{\bf Q}}_{\breve{\bf L},N}^{\epsilon}{\bf f}$ and their approximation errors.
@@ 5 Conclusions and remarks
  • We present a new approach for calculating Fourier extensions, which can effectively overcome the issue of oscillations in the extended region. This idea can also be applied to other types of frame approximations.
@@
  • Existing Fourier extensions consider the overall approximation of the function, and the singularity of the function at any position can substantially impact the approximation result. The results of this paper and [21] actually point out the potential possibility of piecewise approximation, which will be the focus of our further work.

### v3

@@ 3.3 The influence of the parameter $T$ on calculation accuracy
In actual calculations, (29) tells us that when approximating functions of the same frequency, an increase in $T$ means that a larger $N$ is required, and a larger $\gamma$ corresponds to a larger $M$, which will affect the final calculation efficiency. This is not the focus of this paper. For a fast algorithm based on boundary intervals data, please refer to zhao2024fast .
@@ 4 Numerical tests
We show the approximation results of the following six test functions:
$f_{1}(t)=t^{2},\quad f_{2}(t)=\exp(t),\quad f_{3}(t)=e^{\sin(2.7\pi t)+\cos(\pi t)},$
$f_{4}(t)=\text{erf}(10t),\quad f_{5}(t)=\cos(10t^{2}),\quad f_{6}(t)=\cos\left(\frac{10}{1+25t^{2}}\right).$
Fig. 5 shows the weighted projection ${{\bf Q}}_{\breve{\bf L},N}^{\epsilon}{\bf f}$ of $f_{i}$, $i=1,2,\ldots 6$. From the previous analysis, we know that $p=0$ corresponds to the original method, so it shows the characteristic of oscillation around 0 in the extension region. When $p$ takes other values, the function has different forms in the extension region, which is related to the properties of the approximated function. It can be seen that the proposed method effectively alleviates the high-frequency oscillation of the original method in the extension region.
(a) $f_{1}(t)$
(b) $f_{2}(t)$
(c) $f_{3}(t)$
(d) $f_{4}(t)$
(e) $f_{5}(t)$
(f) $f_{6}(t)$
Figure 5: Calculation results of ${{\bf Q}}_{\breve{\bf L},N}^{\epsilon}{\bf f}$ for the various test functions ($N=200$).
Figs. 6-8 show how the approximation errors of different test functions and their first-order and second-order derivatives vary with $N$. The derivatives are calculated by trigonometric interpolation using the FFT on $[-T,T]$. It can be seen that the approximation errors of the functions are very close for different values of $p$, and the proposed method has a significant improvement on the errors of the first-order and second-order derivatives. We also noticed that after $N$ reaches a certain threshold, the error of the derivative calculation tends to get worse as $N$ increases, which is caused by the method we use to calculate the derivatives.
(a) $f_{1}(t)$
(b) $f_{2}(t)$
(c) $f_{3}(t)$.
(d) $f_{4}(t)$
(e) $f_{5}(t)$
(f) $f_{6}(t)$
Figure 6: Approximation errors for the various test functions.
(a) $f_{1}(t)$
(b) $f_{2}(t)$
(c) $f_{3}(t)$
(d) $f_{4}(t)$
(e) $f_{5}(t)$
(f) $f_{6}(t)$
Figure 7: Approximation error of the first-order derivative for various test functions.
(a) $f_{1}(t)$
(b) $f_{2}(t)$
(c) $f_{3}(t)$
(d) $f_{4}(t)$
(e) $f_{5}(t)$
(f) $f_{6}(t)$
Figure 8: Approximation error of the second-order derivative for various test functions.
@@ 5 Conclusions and remarks
  • We present a new approach for calculating Fourier extensions, which can effectively mitigate the issue of oscillations in the extended region. This idea can also be applied to other types of frame approximations.
@@
  • Existing Fourier extensions consider the overall approximation of the function, and the singularity of the function at any position can substantially impact the approximation result. The results of this paper and zhao2024fast actually point out the potential possibility of piecewise approximation, which will be the focus of our further work.
Data Availability
Derived data supporting the findings of this study are available from the authors on request.